OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 1) and (-2, 13)? Answer 4x + y = 5 4x - y = 5 x + 4y = 5 x - 4y = 5

OpenStudy (cwrw238):

the slope is (13-1) / (-2-1) = -4 so the equation is of the form y = -4x + b - that should narrow the choice

OpenStudy (cwrw238):

also x=1 , y=1 will satisfy the equation

OpenStudy (cwrw238):

so value of b can be found 1 = -4 + b b = 5 y = -4x + 5 is correct rearrange this to find the correct one in your list

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

umm one sec let me try

OpenStudy (anonymous):

4x + y = 5 ?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

lol ok thanks

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

u could have solved the by plugging in x=1, y=1 and x = -2 y=13 into each equation to see if they fit but thats a bit tedious

OpenStudy (cwrw238):

- especially if the correct one was the last

OpenStudy (anonymous):

Ok thanks for your help

OpenStudy (cwrw238):

yw